OpenStudy (anonymous):

Find the equation of the line that is parallel to the line formed by the equation x + 3y = -3 and passes through the point (-18, -1).

OpenStudy (anonymous):

The slope of the line is -1/3 and line must satisfy (-18,-1) so y = -1/3 (x - (-18)) -1 y = -1/3 x -7